### v1

# Evaluating, Synthesizing, and Enhancing for Customer Support Conversation

Abstract
Effective customer support requires not only accurate problem-solving but also structured and empathetic communication aligned with professional standards. However, existing dialogue datasets often lack strategic guidance, and real-world service data is difficult to access and annotate. To address this, we introduce the task of Customer Support Conversation (CSC), aimed at training customer service supporters to respond using well-defined support strategies. We propose a structured CSC framework grounded in COPC guidelines, defining five conversational stages and twelve strategies to guide high-quality interactions. Based on this, we construct CSConv, an evaluation dataset of 1,855 real-world customer–agent conversations rewritten using LLMs to reflect deliberate strategy use, and annotated accordingly. Additionally, we develop a role-playing approach that simulates strategy-rich conversations using LLM-powered roles aligned with the CSC framework, resulting in the training dataset RoleCS. Experiments show that fine-tuning strong LLMs on RoleCS significantly improves their ability to generate high-quality, strategy-aligned responses on CSConv. Human evaluations further confirm gains in problem resolution. All code and data will be publicly available at https://github.com/aliyun/qwen-dianjin.
1 Introduction
Figure 1: An example dialogue between a service supporter (left) and a customer (right), showing support strategies (noted in parentheses) used by the supporter. The conversation is organized into five stages of the proposed CSC framework, shown in colored boxes.
Customer support aims to help users resolve product- or service-related issues through effective and context-aware communication. While large-scale dialogue systems have received growing attention in recent years (Budzianowski et al. 2018; Rashkin et al. 2019; Peskov et al. 2019; Liu et al. 2021), customer support conversations remain underexplored in the NLP community, [^1] This paper uses dialogue and conversation interchangeably. largely due to the scarcity of publicly available benchmarks and the sensitive, domain-specific nature of support interactions. Meanwhile, large language models (LLMs) like GPT-3 (Brown et al. 2020) have shown impressive capabilities in open-domain dialogue generation (Zhang et al. 2020b; Bae et al. 2022; Thoppilan et al. 2022), but their ability to generate realistic and effective customer support conversations remains underexamined.
Customer support guidelines from COPC, the internationally recognized standard for customer experience management, [^2] https://www.copc.com/copc-standards/ emphasize that high-quality support often depends on structured communication strategies, similar to those used in emotional support settings (Liu et al. 2021). As shown in Figure 1, [^3] The example is translated into English for better readability. the service supporter begins with the Greeting strategy to initiate the conversation and explore the customer’s issue. Upon recognizing the customer’s negative emotion, the supporter adopts the Emotional Management strategy, expressing empathy and understanding to alleviate the customer’s distress. After understanding the problem, the supporter employs the Information Delivery strategy to provide clear and actionable guidance, followed by the Feedback Request strategy to check for further concerns. Finally, the conversation ends with the Appreciation and Closure strategy, ensuring a respectful and reassuring followed by Feedback Request to check for further concerns.
Despite the importance of effective customer support, research on real-time customer service conversations remains limited, mainly due to a lack of task-specific design and high-quality annotated data. Many studies (Xu et al. 2017; Oraby et al. 2017; Cui et al. 2017; Mesquita, Martins, and Almeida 2022) rely on asynchronous, exchanges (e.g., Twitter), where interactions span minutes to days. These settings differ significantly from the immediate, uninterrupted nature of real-time support. Moreover, effective support requires not only resolving issues but also showing empathy and emotional support. Yet, most task-oriented dialogue datasets (Wen et al. 2017; Budzianowski et al. 2018; Peskov et al. 2019; Gung et al. 2023) lack the intentional use of supportive strategies like emotional management or empathetic closure, which are vital for high-quality customer service.
To address this gap, we introduce the task of customer support conversation (CSC), to facilitate the training of customer service supporters. The goal is to help them respond with appropriate strategies that combine accurate solutions with empathetic communication. Based on COPC standards and inspired by the emotional support framework (Liu et al. 2021), we develop a CSC-specific framework with five dialogue stages and twelve support strategies. Using it, we build CSConv, a high-quality dataset adapted from real service interactions and refined for structured strategy use. To address the lack of high-quality training data, we also develop a role-playing approach that creates RoleCS, a synthetic dataset of strategy-rich conversations by assigning LLMs distinct roles aligned with the CSC framework. Fine-tuning on RoleCS significantly boosts LLMs’ ability to generate strategy-aligned and effective responses on CSConv.
2 Related Work
2.1 Task-Oriented Conversation Datasets.
The goal of CSConv, which is to generate conversations based on real spoken customer service interactions, differs significantly from previous task-oriented dialogue datasets. Many earlier synthetic datasets have been created using the Wizard of Oz (WOZ) framework (Kelley 1984), where one person acts as the system and another as the user. Wen et al. (2017) introduce a crowdsourced version of the WOZ framework to collect domain-specific dialogue data more efficiently. Other task-oriented dialogue datasets include Frames (El Asri et al. 2017), MultiWOZ (Budzianowski et al. 2018), MultiDoGO (Peskov et al. 2019), EMPATHETICDIALOGUES (Rashkin et al. 2019), Taskmaster-1 (Byrne et al. 2019), ESConv (Liu et al. 2021), and NATCS (Gung et al. 2023), each of which explores different domains and annotation schemes for developing conversational agents.
Among them, ESConv is closely related to our work, as both it and CSConv aim to enhance service quality using structured support strategies. However, their construction methods differ significantly: ESConv uses a crowdsourced WOZ setup, while CSConv rewrites real-world service dialogues using high-performing LLMs.
2.2 Role-Playing in Conversation Generation
Recent advances in LLMs have enabled the use of role-playing agents for conversation generation. For example, Bae et al. (2022) generate dialogues aligned with specific roles, while Yang et al. (2024b) simulate characters and settings to produce coherent interactions. Wu et al. (2024) create interactive narrative dramas, and Ye et al. (2025) focus on emotional support conversations. Building on the Ye et al. (2025), we design role-playing agents for customer support with clearer role definitions and responsibilities to more closely reflect real-world service interactions.
Figure 2: Overview of the CSC framework’s five stages, each paired with recommended support strategies (see Table 9 in Appendix A). The typical flow is: ① Connecting → ② Identifying → ③ Exploring → ④ Resolving → ⑤ Maintaining (black arrows), but it can be adjusted based on the specifics of each conversation (dashed arrows).
3 CSC: Customer Support Conversation
In customer support conversation (CSC), the service supporter aims to accurately identify customer issues while also addressing emotional needs. This dual focus enables effective solutions that enhance customer satisfaction and improve problem-resolution efficiency. The supporter is considered effective when both the issue and emotional concerns are properly addressed. To build an evaluation dataset for CSC, we first introduce the CSC framework in Section 3.1, followed by the data construction process in Section 3.2 and dataset statistics in Section 3.3. Finally, we define the CSC task in Section 3.4.
3.1 CSC Framework
The CSC framework organizes the customer support process into five stages, each with recommended strategies designed to enhance the quality and effectiveness of interaction.
Stages.
Building on the three core stages of supportive communication from Hill (2019) and the COPC practical guidelines, we work with domain experts to refine and extend this structure for customer support. The resulting CSC framework defines five stages: Connecting (greeting and rapport building), Identifying (understanding the customer’s issue and emotional state), Exploring (discussing and evaluating potential solutions), Resolving (delivering and confirming resolution), and Maintaining (closing the interaction while preserving the customer relationship). Figure 2 illustrates the flow across these stages.
Importantly, these stages are not rigid steps but modular components that can appear in various orders or combinations depending on the nature of the conversation. For example, even when no solution is reached, a supporter may still provide empathy (Exploring), acknowledge service limitations (Resolving), and close the conversation professionally (Maintaining). This flexible structure enables consistent analysis and strategy modeling across both successful and challenging customer support scenarios.
Strategies.
In parallel, we work with domain experts to define twelve actionable support strategies aligned with the five stages: Greeting (GT), Identity Verification (IV), Emotional Management (EM), Restatement or Paraphrasing (RP), Problem Refinement (PR), Providing Suggestions (PS), Information Delivery (ID), Resolution Implementation (RI), Feedback Request (FR), Appreciation and Closure (AC), Relationship Continuation (RC), and Others. These strategies provide practical guidance for managing both task-related issues and emotional support throughout the customer interaction. For the details of stages and strategies, please refer to Appendix A.
3.2 Dataset Construction
We construct the CSConv dataset using Chinese customer service dialogues collected from our pre-sales and after-sales customer service centers. Prior to our access, all conversations were professionally transcribed, manually corrected, and fully de-identified to ensure privacy protection and integrity. While these raw conversations authentically reflect real-world customer interactions, they often lack consistent structure, making it difficult to systematically annotate support strategies according to the COPC-informed CSC framework. To address this, we employ an LLM to rewrite the conversations while preserving the original semantics and user intent. This controlled rewriting aligns conversations with the defined dialogue stages and strategies, enabling more consistent and interpretable annotations without sacrificing the complexity of real user queries.
Importantly, the goal of constructing CSConv is not to evaluate real-time chatbot performance, but to facilitate the training of customer service supporters by helping them learn to respond using appropriate strategies guided by the COPC framework. In total, we collect 690K conversations spanning eight in-domain topics. To ensure quality, we use a four-stage pipeline to guide data selection and refinement.
  1. 1.
    Pre-filtering: We first apply rule-based filtering to remove low-quality conversations. These rules exclude conversations that are too short or too long, contain overly lengthy utterances, show an imbalance between customer and agent turns, or have a high proportion of ineffective customer responses. Additionally, we use an LLM to exclude conversations with offensive or unprofessional content.
  2. 2.
    Sampling and Rewriting: Up to 500 filtered conversations per topic are sampled and rewritten by an LLM to align with the CSC framework. During this process, the LLM analyzes the original scenario and generates a new conversation that preserves the core issue while improving clarity, structure, and emotional engagement. For each agent turn, the LLM selects an appropriate support strategy based on conversation context, occasionally using Others to maintain conversational naturalness. Customer responses are also refined for coherent interaction.
  3. 3.
    Post-filtering: After rewriting, a second round of rule-based and LLM-based checks verifies structure (e.g., strategy coverage, speaker alternation, and length) and filters out conversations lacking coherence, empathy, or strategic alignment.
  4. 4.
    Manually annotation: Finally, experts certified in COPC review the remaining conversations, evaluating the support agent’s responses for realism, empathy, and adherence to the CSC framework. This results in a curated evaluation set of 1,855 high-quality conversations.
Appendix B provides details on filtering rules, prompt templates, and annotation guidelines. For rewriting, we use DeepSeek-R1 (Guo et al. 2025) instead of GPT-4o (OpenAI 2024), as the latter tends to produce shorter, less emotionally rich dialogues. See Appendix B.5 for a comparison.
3.3 Dataset Analysis
Overview Statistics.
Table 1 compares statistics of CSConv before and after rewriting, covering 1,855 conversations. Rewritten dialogues are longer, averaging 27.27 versus 19.06 utterances, with supporter responses increasing from 41.16 to 48.72 words and customer responses decreasing from 21.60 to 17.17. Notably, strategy use (excluding Others) rises from 55.28% to 97.82%, [^4] We use Qwen2.5-72B-Instruct to assign a support strategy to each supporter response in the dialogues prior to rewriting. indicating more deliberate and guided support. These shifts reflect the goal of rewriting, which is to facilitate the training of customer service supporters in applying appropriate strategies aligned with the COPC framework.
Topic Distribution.
Table 2 presents the distribution of conversations across eight topic. Each topic, excluding Others, accounts for roughly 11% to 16% of the dataset.
- - Number Original Rewritten Total Conversations 1,855 1,855 Utterances 35,350 50,587 Avg. Utterance Number 19.06 27.27 Avg. Utterance Length 31.48 33.27 Supporter Utterances 17,862 25,810 Avg. Utterance Number 9.63 13.91 Avg. Utterance Length 41.16 48.72 Strategy Use (w/o Others) 55.28% 97.82% Customer Utterances 17,488 24,777 Avg. Utterance Number 9.43 13.36 Avg. Utterance Length 21.60 17.17
Table 1: Statistics of CSConv before and after rewriting.
Topic Num Proportion Account and Transaction Management 265 14.3% Product Consultation 242 13.0% Technical Support and Online Services 295 15.9% Complaints and Dispute Resolution 263 14.2% Marketing and Promotion Activities 211 11.4% Risk Management and Security 266 14.3% Financial Consulting and Planning 263 14.2% Others 50 2.7% Overall 1,855 100.0%
Table 2: Distribution of topics in CSConv.
Figure 3: Strategy proportion of CSConv.
Strategy Distribution.
As shown in Figure 3, the most commonly used strategies are Information Delivery (14.9%), Emotional Management (11.9%), and Provide Suggestions (10.0%). This highlights the dual importance of delivering clear information and managing customer emotions in effective support conversations.
Additional statistics on strategy transitions and unique strategy distributions are provided in Appendix C.
3.4 CSC Task Definition
We denote a customer support conversation as $D=\{\left(P_{i},T_{i},U_{i}\right)\}_{i=1}^{N}$, consisting of $N$ turns exchanged between a supporter $S$ and a customer $C$. For each turn, $P_{i}\in\{S,C\}$ denotes the speaker, $U_{i}$ is the utterance text, and $T_{i}$ represents the response strategy used. Strategies are selected from a predefined set $G$, and are assigned only to the supporter’s turns. That is, if $P_{i}=C$, then $T_{i}=\text{NULL}$.
Following the task of emotional support conversation (Liu et al. 2021; Ye et al. 2025), we define the CSC task as generating the supporter’s response. Specifically, at turn $k$, where $P_{k}=S$, the model receives the conversation history $X_{k}=\{\left(P_{i},T_{i},U_{i}\right)\}_{i=1}^{k-1}$ as input. The CSC task is then divided into two subtasks: [^5] We further investigate how strategy prediction contributes to improving response generation in Section 6.3.
  1. 1.
    Strategy Prediction: Predict the appropriate support strategy $T_{k}\in G$ based on the conversation history $X_{k}$.
  2. 2.
    Response Generation: Generate a response $U_{k}$ conditioned on both the predicted strategy $T_{k}$ and the conversation history $X_{k}$, ensuring that the output is consistent with both the customer’s needs and the strategic intent.
4 Synthetic Conversation Generation with Role-Playing Agents
Figure 4: Illustration of synthetic conversation generation using role-playing agents.
Fine-tuning LLMs has shown effective for improving performance in task-specific dialogue applications. However, building high-quality, multi-turn customer support datasets remains challenging due to the need for domain expertise, contextual coherence, and dialogue diversity. This limits both the scalability and scenario coverage. While recent studies use LLMs for dataset augmentation via rewriting or imitation, they often yield limited variation in dialogue flow and support strategy. To address this, inspired by Ye et al. (2025), we adopt a multi-role role-playing framework to generate synthetic customer support dialogues that are diverse, coherent, and representative of real-world scenarios.
4.1 Role Playing Conversation Generation
As shown in Figure 4, our role-playing framework involves five roles: Planner, Supporter Assistant, Supporter, Customer Assistant, and Customer. Each role serves a distinct purpose. The Planner defines the dialogue scenario and sets the Customer’s communication goal. The main interaction occurs between the Supporter and the Customer, simulating a real customer service exchange. Meanwhile, their respective Assistants, the Supporter Assistant and the Customer Assistant, offer strategic guidance to help them fulfill their roles more effectively. We use deepseek-r1 (Guo et al. 2025) to simulate all roles, with detailed prompts available in Appendix E. We describe each role below.
Planner
Before the conversation begins, the Planner selects a topic $e$ from a predefined list $E$ of customer topics in Table 2 (excluding Others). It then samples a customer profile $o$ from a character pool $O$ that contains diverse customer personas. Using the selected topic $e$ and customer profile $o$, the Planner prompts an LLM $\mathcal{M}$ to generate a detailed service scenario $e^{\prime}$ along with a corresponding communication goal $g$, denoted as $(g,e^{\prime})=\mathcal{M}\left(o,e\right)$. This setup ensures that each dialogue is grounded in a realistic context.
Supporter Assistant
The Supporter Assistant recommends an appropriate support strategy $t$ from the predefined strategy set $G$, based on the current dialogue history $h_{s}$ and the scenario $e^{\prime}$. The recommendation is generated by querying the LLM $\mathcal{M}$, i.e., $t=\mathcal{M}(h_{s},G,e^{\prime})$. This strategic guidance helps the Supporter stay aligned with the customer’s needs and maintain coherent throughout the conversation.
Supporter
Guided by the support strategy $t$ suggested by the Supporter Assistant, the Supporter generates a contextually appropriate response $r_{s}$. This response is produced by the LLM $\mathcal{M}$, conditioned on the dialogue history $h_{s}$, the strategy $t$ and the scenario $e^{\prime}$, i.e., $r_{s}=\mathcal{M}(h_{s},t,e^{\prime})$.
Customer Assistant
The Customer Assistant guides the conversation by generating the next direction $d$, ensuring alignment with the customer’s communication goal $g$. To do so, it prompts the LLM $\mathcal{M}$ using the current dialogue history $h_{c}$, the customer’s goal $g$, and the scenario $e^{\prime}$. This helps the customer maintain coherent and goal-oriented behavior throughout the conversation. Formally, the direction is generated as $d=\mathcal{M}(h_{c},g,e^{\prime})$.
Customer
The Customer generates a response $r_{c}$ based on the current dialogue history $h_{c}$, the intended conversational direction $d$, the character profile $o$, and the scenario $e^{\prime}$. This ensures that the response is both contextually appropriate and consistent with the customer’s persona. Formally, the generation process is denoted as: $r_{c}=\mathcal{M}(h_{c},d,o,e^{\prime})$.
Diversity in conversations relies heavily on the richness of customer personas (Wang et al. 2025). To support this, we construct a rich character profile pool $O$ that captures a wide spectrum of customer backgrounds, behaviors, and communication styles, thereby enhancing the realism and variability of the simulated interactions.
Character Profile Pool.
We design a comprehensive character profile template specifically for customer personas, incorporating attributes such as demographics, financial status, and communication preference. To populate this template, we use Qwen2.5-72B-Instruct to automatically extract and complete profile information from 15,980 real-world customer service dialogues. To reduce redundancy and ensure profile diversity, each structured profile is converted into a free-text description, and pairwise cosine similarity is computed using Qwen’s text-embedding-v2. Profiles exceeding a similarity threshold of 0.85 are considered redundant and removed. Following this filtering process, we retain 1,948 distinct customer profiles in our final pool. Additional details, including the profile template and construction prompts, are provided in Appendix D.
4.2 Synthetic Dataset
The Planner generates one conversation per unique pair of customer topic and profile. With (|E|1)(|E|-1) topics (excluding Others) and $|O|$ profiles, this yields 13,636 conversations. After applying quality filters detailed in Appendix F, we retain 11,232 high-quality conversations, forming the final training (or fine-tuning) dataset, RoleCS. To assess whether customer utterances reflect assigned profiles, we randomly sample 500 dialogues and measure word overlap between customer utterances and their corresponding aligned profiles, using random profiles for comparison. As shown in Figure 5, word overlap with aligned profiles rises rapidly at first as profile information is elicited, then increases more gradually once most information has been mentioned. Throughout, word overlap with aligned profiles remains consistently higher than with random profiles, indicating that customer utterances naturally reflect profile details. See Appendix G for more detailed analysis of RoleCS.
Figure 5: Word overlap between cumulative customer utterances and profile (Aligned vs. Random).
5 Experimentation
Model Size B-2 B-4 R-L BS BR ACC B-2 B-4 R-L BS BR ACC
Evaluation with reference context Evaluation with generated context
GPT-4o - 8.13 2.97 4.12 64.26 51.27 42.58 6.41 2.07 2.22 62.95 51.48 36.29
DeepSeek-R1 671B 11.67 5.09 8.44 66.57 52.09 39.78 8.41 3.11 6.27 64.80 52.01 35.23
DeepSeek-V3 671B 11.57 4.95 7.04 66.09 53.10 41.99 8.43 3.33 4.14 64.13 53.09 36.54
LLaMA3.1-Instruct 8B 4.28 1.44 3.62 58.68 38.84 17.16 2.62 0.89 1.31 55.06 35.97 13.75
    + RoleCS 8B 11.06 4.77 6.93 66.21 52.13 42.15 8.80 3.72 4.52 64.57 49.85 35.73
LLaMA3.1-Instruct 70B 6.85 2.38 4.10 63.14 47.53 38.78 5.57 1.76 2.59 62.24 46.83 30.36
    + RoleCS 70B 11.73 5.11 7.64 66.62 53.69 42.79 9.62 4.00 4.89 65.15 50.91 39.44
Qwen2.5-Instruct 7B 6.73 2.29 3.80 62.75 48.21 18.78 5.11 1.53 2.43 61.62 47.43 19.39
    + RoleCS 7B 11.23 4.80 7.39 66.35 52.14 42.96 8.61 3.51 4.13 64.42 48.80 30.52
Qwen2.5-Instruct 72B 8.61 3.23 5.41 64.63 52.49 37.22 6.28 2.00 3.59 63.14 50.55 30.93
    + RoleCS 72B 12.15 5.32 7.97 66.85 54.49 43.29 9.38 3.90 4.35 64.89 53.65 36.02
Table 3: Performance comparison on CSConv. The best and second best results are in bold and underlined, respectively.
Test-Time Prompting.
For both subtasks outlined in Section 3.4, we adopt a unified single-prompt method. Given the conversation history, the LLM, whether fine-tuned or not, is prompted once to first identify the appropriate support strategy and then generate the corresponding response. The full prompt is provided in the Appendix I.
5.1 Experimental Settings
Models.
We evaluate several widely used LLMs on the CSC task, including GPT-4o (OpenAI 2024), DeepSeek-R1 (Guo et al. 2025), DeepSeek-V3 (Liu et al. 2024), Qwen-2.5-7B/72B-Instruct (Yang et al. 2024a), and LLaMA-3.1-8B/70B-Instruct (Grattafiori et al. 2024). To evaluate the impact of the proposed RoleCS, we fine-tune the Qwen and LLaMA models using 137,406 fine-tuning instances extracted from RoleCS, formatted consistently with our single-prompt in test-time. Fine-tuning and inference configurations are provided in Appendix J.
Metrics.
Following prior work (Liu et al. 2021; Ye et al. 2025), we adopt a diverse set of metrics to assess the response quality. These include BLEU-n (B-n(Papineni et al. 2002) and ROUGE-L (R-L(Lin 2004) for measuring lexical overlap, BERTScore (BS(Zhang et al. 2020a) and BLEURT (BR(Sellam, Das, and Parikh 2020) for semantic similarity. In addition, to evaluate alignment with the intended support strategy, which is a core aspect of the CSC task, we report strategy prediction accuracy (ACC).
5.2 Experimental Results
Table 3 presents the main results on CSConv under two evaluation settings: (1) evaluation with reference context, which all LLMs are evaluated using the same gold history for fair comparison, and (2) evaluation with generated context, which assesses performance when relying on model-generated histories, thus reflecting the ability to maintain coherence and relevance without ground truth context. From the results, we have the following observations:
  • Larger models tend to perform better among non–fine-tuned LLMs. Additionally, Chinese-centric models like Qwen and DeepSeek outperform more general models such as LLaMA and GPT, indicating that alignment with language and cultural context benefits CSC performance.
  • RoleCS proves highly effective, as fine-tuning on it significantly improve performance across all metrics. In particular, Qwen2.5-Instruct-72B, after fine-tuning, matches or surpasses DeepSeek-R1, a strong baseline for Chinese-language tasks.
  • Evaluation with generated context shows similar performance trends as with reference context, but with lower absolute scores. This drop reflects the difficulty of maintaining consistency and quality in multi-turn conversations when relying on model-generated dialogue history, echoing findings in prior work such as Ye et al. (2025).
6 Discussion
We conduct further analysis under the reference context setting to examine the effectiveness of our role-playing approach in synthetic conversation generation.
6.1 Impact of Role-Playing on Data Quality
To evaluate our role-playing framework, we compare RoleCS with two baselines:
  • Baseline 1: Conversations are generated via in-context learning without any role-playing.
  • Baseline 2: Role-playing is applied, but without the Supporter Assistant agent.
All three datasets are generated using Deepseek-R1 with equal conversation counts. For fairness, we fine-tune Qwen2.5-7b-Instruct on each. As shown in Table 4, role-playing improves performance over in-context learning (Baseline 2 vs. Baseline 1), and adding the Supporter Assistant further boosts results (RoleCS vs. Baseline 2). These results highlight the value of strategy-rich training data and the importance of high-quality dataset construction.
SFT Data B-2 B-4 R-L BS BR ACC
Baseline 1 8.51 3.09 5.35 64.35 46.70 33.09
Baseline 2 9.40 3.26 6.98 65.22 50.18 36.17
RoleCS 11.23 4.80 7.39 66.35 52.14 42.96
Table 4: Performance comparison of fine-tuning on different datasets.
6.2 Effect of Synthetic Dataset Size
To understand how dataset size influences performance, we split the RoleCS dataset into subsets of {0, 3K, 6K, 9K, All} dialogues. We fine-tune Qwen2.5-7B-Instruct on each subset individually. As shown in Figure 6, the most significant gains occur with the initial 3K examples. Beyond this point, performance improvements become marginal, particularly for ROUGE-L. These results suggest that even a modest amount of high-quality synthetic data can bring strong gains, with limited benefits from adding more.
Figure 6: Performance comparison under different synthetic conversation dataset sizes.
6.3 Impact of Support Strategy Guidance
Strategy B-4 BS ACC
Vanilla 3.11 64.43 -
Predict 3.23 64.63 37.22
Oracle 3.80 66.34 100.00
Table 5: Performance comparison across different support strategy variants.
We evaluate the effect of incorporating support strategies using three variants of Qwen2.5-72B-Instruct without fine-tuning: (1) Vanilla, where the model generates a response without any support strategy guidance; (2) Predict, our default setup, where the model first predicts the strategy and then generates a response; and (3) Oracle, where the ground-truth strategy is provided in the prompt.
As shown in Table 5, the Predict variant slightly outperforms the Vanilla, indicating that even simple strategy prediction improves response quality. The Oracle variant performs best, highlighting that more accurate strategy prediction can further enhance the performance of CSC task.
6.4 Evaluation with LLMs and Human as Judges
[…]
Table 6: Model performance evaluated by GPT-4o, Qwen-Plus, and human judges. $\Delta$ denotes fine-tuned models.
To mitigate potential bias (e.g., GPT-4o favoring its own outputs), we use both GPT-4o and Qwen-Plus [^6] https://help.aliyun.com/zh/model-studio/models to evaluate response quality across six dimensions: accuracy, helpfulness, understanding, coherence, informativeness, and empathy, each scored on a 0–100 scale. We report the average overall scores from both GPT-4o-Judge and Qwen-Plus-Judge, using the evaluation prompt in Appendix K. As shown in Table 6, both judges reveal consistent performance patterns. Notably, fine-tuned Qwen2.5-Instruct-72B and LLaMA3.1-Instruct-70B outperform other models, even exceeding DeepSeek-R1 in overall quality.
Agreement Kappa
GPT-4o-Judge & Human Annotators 0.658
Human Annotators 0.628
Table 7: Fleiss’ Kappa scores.
For human evaluation, we randomly select 100 conversations and have them independently rated by three professional annotators on a 1-5 Likert scale (Joshi et al. 2015) across the same six dimensions. Table 6 shows the average overall scores, which align with trends from the LLM-based evaluations. Table 7 reports Fleiss’ Kappa (Fleiss 1971) scores among annotators and between GPT-4o-Judge and the annotators. The results show strong inter‑rater agreement and confirm that GPT-4o-Judge aligns well with human judgment.
7 Conclusion
This paper addresses the challenges of customer support conversations (CSCs) in the NLP field by introducing CSConv, a high-quality dataset grounded in support strategies, and a role-playing framework for generating realistic, goal-driven dialogues. Our approach significantly enhances LLMs’ ability to generate coherent, context-aware, and empathetic responses in customer service scenarios. We believe that our contributions will inspire further research in this field, promoting advancements in empathetic and effective customer support technologies.
Appendix
Appendix A Stages and Strategies in CSC Framework
Table 8 presents the five stages in the CSC framework while Table 9 shows the twelve strategies in the framework.
Stage Description
Connecting Greeting and establishing connection
Identifying Understanding and identifying problems
Exploring Seeking solution
Resolving Resolving and confirming
Maintaining Ending and maintaining relationship
Table 8: Five stages in the CSC framework.
Strategy Stages Description
Greeting (GT) Utilize friendly and professional language to greet customers, creating a warm communication atmosphere.
Identity Verification (IV) Ensure the accuracy and security of the service by asking for the customer’s basic information.
Emotional Management (EM) Express understanding and care for the customer’s feelings to help alleviate negative emotions.
Restatement or Paraphrasing (RP) Restate the customer’s issue to ensure accurate understanding.
Problem Refinement (PR) Employ detailed inquiries to fully and accurately comprehend customer needs.
Providing Suggestions (PS) Offer professional advice or action steps based on the customer’s issue.
Information Delivery (ID) Clearly explain relevant company policies, processes, or steps to help customers understand the basis of solutions.
Resolution Implementation (RI) Execute the agreed-upon solution, ensuring all steps are followed as planned, and update the customer on the progress.
Feedback Request (FR) Seek customer feedback after the issue has been addressed to gauge their satisfaction and identify potential areas for improvement.
Appreciation and Closure (AC) End the conversation positively, ensuring the customer feels valued and laying a solid foundation for future interactions.
Relationship Continuation (RC) Guide customers towards future service or product updates, ensuring they understand how to continue receiving support and service in the future, thereby establishing a bridge for further interaction.
Others Situations that do not belong to the above eleven strategies.
Table 9: Twelve strategies in the CSC framework. The cells of lightblue, lightgreen, lightpurple, lightyellow, and lightpink represent the Connecting, Identifying, Exploring, Resolving, and Maintaining stages, repsectively.
Appendix B More Details of Data Construction
In the following, we present more details to the pre-filtering, rewriting and post-filtering in data construction. Then, we compare written conversations by DeepSeek-R1 and GPT-4o.
B.1 Rules and Prompt for Pre-filtering
To ensure the quality and suitability of the initial customer support conversations, we apply a set of rule-based pre-filtering rules prior to rewriting. These rules help filter out conversations that are structurally inadequate or contain limited user engagement:
  • Rule 1: Dialogue Length Constraint. Each conversation must contain more than 6 but fewer than 60 utterances to ensure sufficient interaction.
  • Rule 2: Utterance Length Constraint. No single utterance should exceed 500 characters.
  • Rule 3: Speaker Balance. The number of utterances from the service supporter must not exceed twice the number of customer utterances, promoting balanced participation.
  • Rule 4: Customer Utterance Effectiveness. At least 70% of customer utterances must be considered effective, where an utterance is defined as effective if it contains more than 3 characters, ensuring active and meaningful user input.
In addition, we utilize an LLM (Qwen2.5-72B-Instruct) to automatically assess whether a conversation is of low quality. A conversation is flagged as low quality if it contains any of the following: (1) explicitly offensive, abusive, or inappropriate language, or (2) clear signs of unprofessional behavior from the service supporter, such as impatience, indifference, or disrespect. Figure 7 shows the prompt used for this evaluation.
Figure 7: Prompt template for pre-filtering.
B.2 Prompt for Rewriting
Figure 8 shows the prompt for rewriting.
Figure 8: Prompt template for rewriting.
B.3 Rules and Prompt for Post-filtering
After the conversations are rewritten, we apply a series of post-filtering rules to ensure quality and structural consistency:
  • Rule 1: Minimum Utterance Requirement. Each conversation must contain at least 10 utterances and at most 50 utterance to ensure sufficient conversational depth.
  • Rule 2: System Message Removal. Any system-generated messages (e.g., “(System Auto-Push)”) at the end of conversations are removed to maintain textual clarity.
  • Rule 3: Strategy Presence Check. The support strategies Greeting, Identity Verification, and Appreciation and Closure must each appear at least once in the dialogue to reflect a complete and well-structured interaction.
  • Rule 4: Speaker Alternation Constraint. The customer and service supporter must alternate speaking in the conversation.
After applying above rules, we further assess the overall quality of the rewritten conversations using an LLM (Qwen2.5-72B-Instruct). The model is prompted to evaluate each dialogue and classify it as either high or low quality based on coherence, naturalness, and adherence to support strategies. Conversations deemed to be of low quality are subsequently discarded. The prompt used for this evaluation is shown in Figure 9.
Figure 9: Prompt template for post-filtering.
B.4 Annotation Guideline
Figure 10 shows the annotation guidelines for customer support conversations, designed by domain experts.
Figure 10: Annotation guideline for evaluating the quality of customer support conversations.
B.5 Example of Rewritten Conversations by Deepseek-R1 and GPT-4o
Figure 11 shows a conversation before rewriting, while Figure 12 and Figure 13 compare the rewritten conversations by GPT-4o and DeepSeek-R1.
Figure 11: An example from real-world customer service conversations.
Figure 12: Rewritten dialogue by GPT-4o.
Figure 13: Rewritten dialogue by DeepSeek-R1.
Appendix C More Statistics of CSConv
Figure 14: Strategies distribution at different conversation progress on CSConv dataset.
Figure 15: Distribution of strategy types across different dialogues on CSConv.
Strategy Transition.
[…]
To analyze strategy transitions, we identify the most frequent 2-hop and 3-hop patterns in the conversations. The results are presented in Table 10. These patterns highlight common strategy flows such as Greeting (GT) $\rightarrow$ Identity Verification (IV) and Information Delivery (ID) $\rightarrow$ Providing Suggestions (PS), reflecting a structured and consistent progression in customer service interactions. The observed transitions align well with standard service protocols, emphasizing coherent dialogue management.
- Pattern Proportion
2-Hop GT $\rightarrow$ IV 5.69%
IV $\rightarrow$ RP 4.84%
ID $\rightarrow$ PS 4.55%
FR $\rightarrow$ AC 4.22%
EM $\rightarrow$ ID 3.59%
3-Hop GT $\rightarrow$ IV $\rightarrow$ RP 4.85%
IV $\rightarrow$ RP $\rightarrow$ EM 2.57%
RC $\rightarrow$ FR $\rightarrow$ AC 2.42%
ID $\rightarrow$ PS $\rightarrow$ RI 1.91%
EM $\rightarrow$ ID $\rightarrow$ PS 1.77%
Table 10: Top-5 2-hop and 3-hop strategy patterns.
Distribution of Strategy Types.
To examine the diversity of support strategies, we analyze the number of distinct strategy types used within individual conversations, as shown in Figure 15. Most conversations employ more than ten different strategy types, demonstrating the broad applicability and flexibility of the predefined strategy set. This variety reflects the dynamic nature of real-world customer service interactions and enhances the overall richness of the dataset.
Appendix D Details of Character Profile Construction
Figure 16: Profile example.
Figure 17: Prompt template for completing a profile based on real-world customer service conversation (In Chinese).
Figure 18: Prompt template for completing a profile based on real-world customer service conversation (In English).
Figure 19: Prompt template for transforming structured profile into descriptive text.
Figure 16 illustrates an example of profiles, which are constructed using the prompt presented in Figure 17 (In Chinese) and Figure 18 (In English). Figure 19 shows the prompt for profile-to-text conversion.
Appendix E Prompts for Generating Synthetic Conversation Dataset
Figure 20: Prompt template used by planner agent.
Figure 21: Prompt template used by supporter assistant agent.
Figure 22: Prompt template used by supporter agent.
Figure 23: Prompt template used by customer assistant agent.
Figure 24: Prompt template used by customer agent.
Figures 20, 21, 22, 23, and 24 illustrate the prompts used by Planner, Supporter Assistant, Supporter, Customer Assistant, and Customer, respectively, during the synthetic dialogue generation process. These prompts guide each role’s behavior to ensure coherence, strategy alignment, and contextual consistency throughout the conversation.
Appendix F Filtering Rules for Synthetic Dialogues
Figure 25: Prompt template used in filtering synthetic conversations.
We first discard dialogues with fewer than 10 or more than 50 utterances, then use the prompt from Figure 25 to filter out low-quality dialogues.
Appendix G Synthetic Dataset Analysis
- - Number
Total Dialogues 11,232
Utterances 263,580
Avg. Utterance Number 23.47
Avg. Utterance Length 57.14
Supporter Utterances 137,406
Avg. Utterance Number 12.23
Avg. Utterance Length 66.98
Customer Utterances 126,174
Avg. Utterance Number 11.23
Avg. Utterance Length 46.43
Table 11: Statistics of the RoleCS dataset.
Figure 26: Strategies distribution at different conversation progress in the RoleCS dataset.
Figure 27: Number of strategy types in a dialogue within the RoleCS dataset.
Similar to the analysis of the CSConv dataset, Table 11 summarizes the statistics of the RoleCS dataset, while Figure 27 shows strategy distribution across conversation progress, and Figure 27 illustrates the variety of strategy types within dialogues.
Appendix H Diversity Analysis
Dataset Distinct-1 Distinct-2 Distinct-3
CSConv 1.09 19.27 46.07
RoleCS 0.95 22.35 55.31
Table 12: Distinct-n scores of 1,000 conversations from CSConv and RoleCS.
Figure 28: Cosine similarity statistics between pairs of distinct dialogues using TF-IDF vectors.
We examine the diversity of the CSConv and RoleCS datasets from two perspectives: lexical diversity and semantic diversity. To assess lexical diversity, we adopt Distinct-n (Li et al. 2016) metric. For a fair comparison, we randomly select 1,000 dialogues from each dataset, removing speaker and strategy information, and utilizing Qwen2.5-7B tokenizer for tokenization. Regarding semantic diversity, we calculate cosine similarity between pairs of distinct dialogues using TF-IDF features (Salton, Wong, and Yang 1975). Table 12 and Figure 28 present the results, respectively.
Appendix I Prompt for Predicting Strategy and Generating Response
Figure 29 presents the prompt for predicting strategy and generating response on the CSC task.
Appendix J Details of Fine-tuning and Inference
For fine-tuning, we use the LLaMA-Factory framework (Zheng et al. 2024), applying LoRA (Hu et al. 2022) with a rank of 8 and a scaling factor of 16. All fine-tunings are conducted on 4 NVIDIA A100 80GB GPUs, with a batch size of 4 and gradient accumulation over 2 steps. The learning rate is set to 3e-5, and fine-tuning is run for 3 epochs to mitigate the risk of overfitting.
For inference, we use the checkpoint from the final epoch and configure the decoding parameters with a top-p value of 0.7 and a temperature of 0.95.
Appendix K Prompt for LLM as Judge
Figure 30 shows the evaluation prompt used when employing GPT-4o and Qwen-Plus as judges.
Figure 29: Prompt template for predicting support strategy and generating response on the CSC task.
Figure 30: Prompt template used by LLM as judge.

### v2

@@
 Customer support aims to help users resolve product- or service-related issues through effective and context-aware communication.
-While large-scale dialogue systems have received growing attention in recent years (Budzianowski et al.
-2018; Rashkin et al.
-2019; Peskov et al.
-2019; Liu et al.
-2021), customer support conversations remain underexplored in the NLP community, [^1] This paper uses dialogue and conversation interchangeably.
+While large-scale dialogue systems have received growing attention in recent years (budzianowski-etal-2018-multiwoz; rashkin-etal-2019-towards; peskov-etal-2019-multi; liu-etal-2021-towards), customer support conversations remain underexplored in the NLP community, [^1] This paper uses dialogue and conversation interchangeably.
 largely due to the scarcity of publicly available benchmarks and the sensitive, domain-specific nature of support interactions.
-Meanwhile, large language models (LLMs) like GPT-3 (Brown et al.
-2020) have shown impressive capabilities in open-domain dialogue generation (Zhang et al.
-2020b; Bae et al.
-2022; Thoppilan et al.
-2022), but their ability to generate realistic and effective customer support conversations remains underexamined.
-Customer support guidelines from COPC, the internationally recognized standard for customer experience management, [^2] https://www.copc.com/copc-standards/ emphasize that high-quality support often depends on structured communication strategies, similar to those used in emotional support settings (Liu et al.
-2021).
+Meanwhile, large language models (LLMs) like GPT-3 (brown-etal-2020-language) have shown impressive capabilities in open-domain dialogue generation (zhang-etal-2020-dialogpt; bae-etal-2022-building; thoppilan-etal-2022-lamda), but their ability to generate realistic and effective customer support conversations remains underexamined.
+Customer support guidelines from COPC, the internationally recognized standard for customer experience management, [^2] https://www.copc.com/copc-standards/ emphasize that high-quality support often depends on structured communication strategies, similar to those used in emotional support settings (liu-etal-2021-towards).
 As shown in Figure 1, [^3] The example is translated into English for better readability.
@@
 Despite the importance of effective customer support, research on real-time customer service conversations remains limited, mainly due to a lack of task-specific design and high-quality annotated data.
-Many studies (Xu et al.
-2017; Oraby et al.
-2017; Cui et al.
-2017; Mesquita, Martins, and Almeida 2022) rely on asynchronous, exchanges (e.g., Twitter), where interactions span minutes to days.
+Many studies (xu-etal-2017-a; oraby-etal-2017-how; cui-etal-2017-superagent; mesquita-etal-2022-dense) rely on asynchronous, exchanges (e.g., Twitter), where interactions span minutes to days.
 These settings differ significantly from the immediate, uninterrupted nature of real-time support.
 Moreover, effective support requires not only resolving issues but also showing empathy and emotional support.
-Yet, most task-oriented dialogue datasets (Wen et al.
-2017; Budzianowski et al.
-2018; Peskov et al.
-2019; Gung et al.
-2023) lack the intentional use of supportive strategies like emotional management or empathetic closure, which are vital for high-quality customer service.
+Yet, most task-oriented dialogue datasets (wen-etal-2017-network; budzianowski-etal-2018-multiwoz; peskov-etal-2019-multi; gung-etal-2023-natcs) lack the intentional use of supportive strategies like emotional management or empathetic closure, which are vital for high-quality customer service.
 To address this gap, we introduce the task of customer support conversation (CSC), to facilitate the training of customer service supporters.
 The goal is to help them respond with appropriate strategies that combine accurate solutions with empathetic communication.
-Based on COPC standards and inspired by the emotional support framework (Liu et al.
-2021), we develop a CSC-specific framework with five dialogue stages and twelve support strategies.
+Based on COPC standards and inspired by the emotional support framework (liu-etal-2021-towards), we develop a CSC-specific framework with five dialogue stages and twelve support strategies.
 Using it, we build CSConv, a high-quality dataset adapted from real service interactions and refined for structured strategy use.
@@
 The goal of CSConv, which is to generate conversations based on real spoken customer service interactions, differs significantly from previous task-oriented dialogue datasets.
-Many earlier synthetic datasets have been created using the Wizard of Oz (WOZ) framework (Kelley 1984), where one person acts as the system and another as the user.
-Wen et al.
-(2017) introduce a crowdsourced version of the WOZ framework to collect domain-specific dialogue data more efficiently.
-Other task-oriented dialogue datasets include Frames (El Asri et al.
-2017), MultiWOZ (Budzianowski et al.
-2018), MultiDoGO (Peskov et al.
-2019), EMPATHETICDIALOGUES (Rashkin et al.
-2019), Taskmaster-1 (Byrne et al.
-2019), ESConv (Liu et al.
-2021), and NATCS (Gung et al.
-2023), each of which explores different domains and annotation schemes for developing conversational agents.
+Many earlier synthetic datasets have been created using the Wizard of Oz (WOZ) framework (kelley-1984-an), where one person acts as the system and another as the user.
+wen-etal-2017-network introduce a crowdsourced version of the WOZ framework to collect domain-specific dialogue data more efficiently.
+Other task-oriented dialogue datasets include Frames (asri-etal-2017-frames), MultiWOZ (budzianowski-etal-2018-multiwoz), MultiDoGO (peskov-etal-2019-multi), EMPATHETICDIALOGUES (rashkin-etal-2019-towards), Taskmaster-1 (byrne-etal-2019-taskmaster), ESConv (liu-etal-2021-towards), and NATCS (gung-etal-2023-natcs), each of which explores different domains and annotation schemes for developing conversational agents.
 Among them, ESConv is closely related to our work, as both it and CSConv aim to enhance service quality using structured support strategies.
@@
 Recent advances in LLMs have enabled the use of role-playing agents for conversation generation.
-For example, Bae et al.
-(2022) generate dialogues aligned with specific roles, while Yang et al.
-(2024b) simulate characters and settings to produce coherent interactions.
-Wu et al.
-(2024) create interactive narrative dramas, and Ye et al.
-(2025) focus on emotional support conversations.
-Building on the Ye et al.
-(2025), we design role-playing agents for customer support with clearer role definitions and responsibilities to more closely reflect real-world service interactions.
+For example, bae-etal-2022-building generate dialogues aligned with specific roles, while yang-etal-2024-crafting simulate characters and settings to produce coherent interactions.
+wu-etal-2024-from create interactive narrative dramas, and ye-etal-2025-sweetiechat focus on emotional support conversations.
+Building on the ye-etal-2025-sweetiechat, we design role-playing agents for customer support with clearer role definitions and responsibilities to more closely reflect real-world service interactions.
 Figure 2: Overview of the CSC framework’s five stages, each paired with recommended support strategies (see Table 9 in Appendix A).
@@
 Stages.
-Building on the three core stages of supportive communication from Hill (2019) and the COPC practical guidelines, we work with domain experts to refine and extend this structure for customer support.
+Building on the three core stages of supportive communication from hill-2019-help and the COPC practical guidelines, we work with domain experts to refine and extend this structure for customer support.
 The resulting CSC framework defines five stages: Connecting (greeting and rapport building), Identifying (understanding the customer’s issue and emotional state), Exploring (discussing and evaluating potential solutions), Resolving (delivering and confirming resolution), and Maintaining (closing the interaction while preserving the customer relationship).
@@
 Appendix B provides details on filtering rules, prompt templates, and annotation guidelines.
-For rewriting, we use DeepSeek-R1 (Guo et al.
-2025) instead of GPT-4o (OpenAI 2024), as the latter tends to produce shorter, less emotionally rich dialogues.
+For rewriting, we use DeepSeek-R1 (guo-etal-2025-deepseek-r1) instead of GPT-4o (openai-2024-gpt4o), as the latter tends to produce shorter, less emotionally rich dialogues.
 See Appendix B.5 for a comparison.
@@
 That is, if $P_{i}=C$, then $T_{i}=\text{NULL}$.
-Following the task of emotional support conversation (Liu et al.
-2021; Ye et al.
-2025), we define the CSC task as generating the supporter’s response.
+Following the task of emotional support conversation (liu-etal-2021-towards; ye-etal-2025-sweetiechat), we define the CSC task as generating the supporter’s response.
 Specifically, at turn $k$, where $P_{k}=S$, the model receives the conversation history $X_{k}=\{\left(P_{i},T_{i},U_{i}\right)\}_{i=1}^{k-1}$ as input.
@@
 While recent studies use LLMs for dataset augmentation via rewriting or imitation, they often yield limited variation in dialogue flow and support strategy.
-To address this, inspired by Ye et al.
-(2025), we adopt a multi-role role-playing framework to generate synthetic customer support dialogues that are diverse, coherent, and representative of real-world scenarios.
+To address this, inspired by ye-etal-2025-sweetiechat, we adopt a multi-role role-playing framework to generate synthetic customer support dialogues that are diverse, coherent, and representative of real-world scenarios.
 4.1 Role Playing Conversation Generation
@@
 Meanwhile, their respective Assistants, the Supporter Assistant and the Customer Assistant, offer strategic guidance to help them fulfill their roles more effectively.
-We use deepseek-r1 (Guo et al.
-2025) to simulate all roles, with detailed prompts available in Appendix E.
+We use deepseek-r1 (guo-etal-2025-deepseek-r1) to simulate all roles, with detailed prompts available in Appendix E.
 We describe each role below.
@@
 Formally, the generation process is denoted as: $r_{c}=\mathcal{M}(h_{c},d,o,e^{\prime})$.
-Diversity in conversations relies heavily on the richness of customer personas (Wang et al.
-2025).
+Diversity in conversations relies heavily on the richness of customer personas (wang-etal-2025-opencharacter).
 To support this, we construct a rich character profile pool $O$ that captures a wide spectrum of customer backgrounds, behaviors, and communication styles, thereby enhancing the realism and variability of the simulated interactions.
@@
 Models.
-We evaluate several widely used LLMs on the CSC task, including GPT-4o (OpenAI 2024), DeepSeek-R1 (Guo et al.
-2025), DeepSeek-V3 (Liu et al.
-2024), Qwen-2.5-7B/72B-Instruct (Yang et al.
-2024a), and LLaMA-3.1-8B/70B-Instruct (Grattafiori et al.
-2024).
+We evaluate several widely used LLMs on the CSC task, including GPT-4o (openai-2024-gpt4o), DeepSeek-R1 (guo-etal-2025-deepseek-r1), DeepSeek-V3 (liu-etal-2024-deepseek-v3), Qwen-2.5-7B/72B-Instruct (yang-etal-2024-qwen2.5), and LLaMA-3.1-8B/70B-Instruct (grattafiori-etal-2024-llama3).
 To evaluate the impact of the proposed RoleCS, we fine-tune the Qwen and LLaMA models using 137,406 fine-tuning instances extracted from RoleCS, formatted consistently with our single-prompt in test-time.
@@
 Metrics.
-Following prior work (Liu et al.
-2021; Ye et al.
-2025), we adopt a diverse set of metrics to assess the response quality.
-These include BLEU-n (B-n(Papineni et al.
-2002) and ROUGE-L (R-L(Lin 2004) for measuring lexical overlap, BERTScore (BS(Zhang et al.
-2020a) and BLEURT (BR(Sellam, Das, and Parikh 2020) for semantic similarity.
+Following prior work (liu-etal-2021-towards; ye-etal-2025-sweetiechat), we adopt a diverse set of metrics to assess the response quality.
+These include BLEU-n (B-n(papineni-etal-2002-bleu) and ROUGE-L (R-L(lin-2004-rouge) for measuring lexical overlap, BERTScore (BS(zhang-etal-2020-bertscore) and BLEURT (BR(sellam-etal-2020-bleurt) for semantic similarity.
 In addition, to evaluate alignment with the intended support strategy, which is a core aspect of the CSC task, we report strategy prediction accuracy (ACC).
@@
   • Evaluation with generated context shows similar performance trends as with reference context, but with lower absolute scores.
-This drop reflects the difficulty of maintaining consistency and quality in multi-turn conversations when relying on model-generated dialogue history, echoing findings in prior work such as Ye et al.
-(2025).
+This drop reflects the difficulty of maintaining consistency and quality in multi-turn conversations when relying on model-generated dialogue history, echoing findings in prior work such as ye-etal-2025-sweetiechat.
 6 Discussion
@@
 Table 7: Fleiss’ Kappa scores.
-For human evaluation, we randomly select 100 conversations and have them independently rated by three professional annotators on a 1-5 Likert scale (Joshi et al.
-2015) across the same six dimensions.
+For human evaluation, we randomly select 100 conversations and have them independently rated by three professional annotators on a 1-5 Likert scale (joshi-2015-likert) across the same six dimensions.
 Table 6 shows the average overall scores, which align with trends from the LLM-based evaluations.
-Table 7 reports Fleiss’ Kappa (Fleiss 1971) scores among annotators and between GPT-4o-Judge and the annotators.
+Table 7 reports Fleiss’ Kappa (fleiss-1971-kappa) scores among annotators and between GPT-4o-Judge and the annotators.
 The results show strong inter‑rater agreement and confirm that GPT-4o-Judge aligns well with human judgment.
@@
 We examine the diversity of the CSConv and RoleCS datasets from two perspectives: lexical diversity and semantic diversity.
-To assess lexical diversity, we adopt Distinct-n (Li et al.
-2016) metric.
+To assess lexical diversity, we adopt Distinct-n (li-etal-2016-diversity) metric.
 For a fair comparison, we randomly select 1,000 dialogues from each dataset, removing speaker and strategy information, and utilizing Qwen2.5-7B tokenizer for tokenization.
-Regarding semantic diversity, we calculate cosine similarity between pairs of distinct dialogues using TF-IDF features (Salton, Wong, and Yang 1975).
+Regarding semantic diversity, we calculate cosine similarity between pairs of distinct dialogues using TF-IDF features (salton-1975-vector).
 Table 12 and Figure 28 present the results, respectively.
@@
 Appendix J Details of Fine-tuning and Inference
-For fine-tuning, we use the LLaMA-Factory framework (Zheng et al.
-2024), applying LoRA (Hu et al.
-2022) with a rank of 8 and a scaling factor of 16.
+For fine-tuning, we use the LLaMA-Factory framework (zheng-etal-2024-llamafactory), applying LoRA (hu-etal-2022-lora) with a rank of 8 and a scaling factor of 16.
 All fine-tunings are conducted on 4 NVIDIA A100 80GB GPUs, with a batch size of 4 and gradient accumulation over 2 steps.